OpenStudy (candycanegone):

Two investments earn an annual income of $621. One investment earns an annual simple interest rate of 8.5%, and the other investment earns an annual simple interest rate of 6.1%. The total amount invested is $9000. How much is invested in each account?

OpenStudy (dumbcow):

\[x + y = 9000\] \[.085x + .061y = 621\] Solve system using substitution \[y = 9000 - x\] \[.085x + .061(9000-x) = 621\] combine like terms \[x(.085-.061) = 621 - 9000(.061)\] \[x = \frac{72}{.024} = 3000\] Therefore \[y = 9000 - 3000 = 6000\]